# Hardwired-Neurons Language Processing Units as General-Purpose Cognitive Substrates

### v1

@@ 3. Design Principles
Figure 3. Hardwired-Neuron architecture. ❶ A conventional cell-embedding neuron contains 2,880 4b multipliers (16 shown) followed by an 8b×2,880 adder tree, where 2,880 is the hidden size in gpt-oss 120 B; ❷ With ME, Hardwired Neurons accept 1b serialized inputs (LSB-first), (1) route the inputs multiplying the same weight value to the same region, (2) perform counting (POPCNT) on these input signals, (3) perform actual multiplication with 16 multipliers (4 shown), (4) sum the results with a 4b×16 adder tree. Note how ❷ is significantly smaller in area than ❶ by reducing the number of multipliers and the strength of adders.
Currently, almost all AI chips are designed as regular grid structures (Figure 2 ❶). Inputs, weights, and outputs are transmitted into/out-of the grid by a regular crossbar of metal wires. By fixing the weights, multipliers could be optimized as multiply-by-constant which is several times simpler in Boolean complexity. For example, FP4 multiply-by-constant is 42.5 Tr on average, $\sim$6× smaller than FP4 multiplier. However, this density level is still far behind the requirement of HNLPU.
In current methodologies, metal layers are only considered during place-and-route (P&R), and the topology of metal wires do not express any specific information. We see this as a waste of resources. The key insight here is that, as the topology of metal wires are three-dimensional, they potentially could embed information in a much higher density than silicon devices. As silicon devices are the area-limiting factors in the design, routing signals through complex metal wires is almost free in both area and energy, compared with logic implemented in standard cells. Please recall that biological neurons in our brains have complex topology of axon-dendrite interconnections in the first place, only then come synaptic weights.
Metal-Embedding Hardwired-Neurons
@@
Then HN carries the multiplying per region, as shown in Figure 3 ❷. As all input signals $x_{1},x_{2},\dots,x_{n}$ in the same region are multiplying with the same constant weight value $a$, instead of performing $ax_{1}+ax_{2}+\dots+ax_{n}$, HN performs $a\left(x_{1}+x_{2}+\dots+x_{n}\right)$ by applying the distributive law. Furthermore, if input signals $x_{1},x_{2},\dots,x_{n}$ are in binary format, it could be serialized from the least-significant bit (LSB) to the most-significant bit (MSB) to further simplify the circuit. After serialization, the arithmetic per region is reduced to a population count (POPCNT) followed by a constant-multiplication. The partial results from regions are then summed with a now much smaller adder tree.
Specifically, a conventional cell-embedding neuron (Figure 3 ❶) requires 2,880 4b×4b multiply-by-constant, followed by an 8b×2,880 adder tree. A ME HN (Figure 3 ❷) reduced the arithmetic strength to (2,880 in total)×1b POPCNT, followed by 16 4b×1b multiply-by-constant and a 4b×16 adder tree. By removing the weight embedding from area-limiting silicon layers, ME HN reduced the area by an order-of-magnitude, which we will demonstrate with post-layout evaluations.
Figure 4. Co-optimizing Metal-Embedding with technology. The embedding metal wires are routed at higher level (M8-M11), leaving other layers (devices, cell-level/module-level interconnections, power and peripherals) homogeneous in structures. As a result, all HNLPU chips could share the same set of photomask at layers other than M8-M11, including all critical layers which require expensive EUV photomask.
Co-optimizing with lithography technology
Another opportunity brought by ME is to co-optimize with lithography to save photomask. Photomask sets account for the dominant part of NRE for HNLPU. Previous chiplet or wafer-scale designs always share one or few sets of photomask across chiplets/wafer-tiles, since they are designed as homogeneous structures. However, each chip in HNLPU is embedding different part of the model weights, they are heterogeneous in nature. Even with significantly reduced area of ME HN, HNLPU still requires 16 full mask sets, each valued around \$ 30 M at 5 nm technology. The total NRE still offset most economical interests considering the typical lifespan of LLMs.
[…]
If critical layers (silicon and lower metals) are made homogeneous, their photomasks could be reused across chips, then the total NRE will be significantly reduced. Therefore, we propose to place the ME only in higher metal layers (M8+) to enable cost-efficient fabrication with DUV, as shown in Figure 4.
This requires a moderate change in the IC design flow. Similar to the sea-of-gates architecture (de Lima and Kinniment, 1995) emerged in 1990s which prefabricate gates on wafers, we prefabricate HNs, resulting in a sea-of-neurons architecture. First, we complete the physical design of a group of homogeneous HNs as a repeating pattern. As homogeneous HN design can no longer assume the number of input signals in each region, the POPCNT is divided into smaller slices and reconfigured with ME, instead of sizing on demand. Specifically, for gpt-oss 120 B there are 106 32b POPCNT slices in each HN, sufficiently reconfigurable for all assignments of the 2,880 inputs. The ports of POPCNT are routed up to M7, left for ME (VIA7) to connect. Then, the layout of HN is copied to fill the major part of die area, equipped with all SoC peripherals, power grid, and clock tree. Finally, each chip is individually wired at M8+ to embed their part of model weights. Figure 5 shows the layout of a randomly selected group of HNs, divided into homogeneous structures (M7-) and ME at the top (M8-M11). ME yielded a congestion-free layout (peak density <<70%, zero overflow) and achieved timing closure at 1 GHz. The final design passed all DRC and LVS sign-off checks.
As a result, all HNLPU chips could share the same set of photomask at all layers other than M8-M11. At the scale of 16 chips, the estimated cost on photomask reduced from \$ 480 M to \$ 65 M. Furthermore, a parameter-only update can share the photomask set or build upon the prefabricated wafers with sea-of-neurons, thus the cost is estimated at \$ 37 M.
Figure 5. Layout of a selected group of HNs.
@@ 4.1. System Integration
The HNLPU system architecture is built upon a 16-module row-column fully-connected fabric. As conceptually illustrated by the logical topology in Figure 6(a), this fabric establishes direct, point-to-point links from each module to all other modules within its row and, simultaneously, to all other modules within its column. This design creates a router-less low-latency network for efficient collective communication patterns (e.g., All-Reduce). Each compute module is a self-contained unit, equipped with a dedicated HBM for storing the KV Cache and the embedding tables.
@@
    For attention score, query-head is all-reduced within the same column groups, while key- and value- heads are reduced to the chip-($\ell\bmod 4$), where $\ell$ denotes the sequence length).
@@ 4.2. Single Chip Architecture
Attention Buffer. On each chip, the Attention Buffer provides 320 MB storage with a high throughput of 64 TB/s. It primarily functions as a KV Cache for the chip’s assigned attention groups. It offloads excess KV entries to HBM only when the on-chip capacity is exceeded. This buffer also stores activation vectors for residual connections in the FFN blocks.
@@ 5. Execution Dataflow
In this section, we introduce the multi-chip interconnect dataflow of HNLPU, covering the model-to-chip mapping and the computing process of a Transformer block. In Section 5.1, we provide an overview of the HNLPU dataflow. Section 5.2 describes the computation of Group-Query Attention. In Section 5.3, we present the data flow for the MoE computation. Section 5.4 details our pipelining strategy and batching inference scheduling.
@@ 5.1. Dataflow Overview
Figure 7 illustrates the dataflow of HNLPU. Our design is driven by the primary goals of distributing computational, KV cache memory access, and thermal loads, while minimizing inter-chip data communication.
@@
The FFN implementation, which uses a Mixture-of-Experts (MoE) architecture, assigns eight experts to each chip, allowing for entirely independent computation with no inter-chip communication. A key exception to our partitioning strategy is the router’s weight matrix, $W_{\text{rout}}$, which is replicated across all chips. This deliberate design trade-off introduces a negligible area overhead—as the router’s weights constitute only about 0.01% of the total model weights—but eliminates the communication latency. Once all transformer blocks have been processed, the final output vector is passed through the unembedding layer, and a new token index is calculated through a sampling operation.
5.2. Grouped-Query Attention
Figure 7.(II) illustrates the computation dataflow of the 64 query heads projection. The token activation $X$ that in the chip array has shape $(1,2880)$. The activation $X$ in all chips is split into four equal slices of $(1,720)$ and each chip takes one of slices to calculate the partial sum of the query tensor. Each chip contains a private, hard-wired slice of $W_{q}$ with shape $(720,1024)$ inside the HN array. Therefore the product is generated locally and without any weight fetch. The four partial products are summed by a column-internal reduce operation, yielding 16-heads query vector $(1,1024)$. The 16 query heads are then reshaped into $(2,8,64)$, representing that in Grouped-Query Attention, every 8 query heads correspond to a single KV head.
Figure 7.(III) follows a similar spatial pattern for the key path. Each chip multiplies its $(1,720)$ input slice with its hard-wired $W_{k}$ slice $(720,128)$ and emits a partial key vector $(1,128)$. After the column-internal reduction and reshape, the new key head of $(2,64)$ is hold in chip# (# $\ell\bmod 4$, where $\ell$ is the token’s position in the sequence). At this point the data layout is intentionally asymmetric: the query vector is fully replicated across columns, whereas each key vector is unique per chip in the sequence dimension, a choice that minimises the traffic of the forthcoming dot-product. The data flow of $X\cdot W_{v}$ is the same as that of $W_{k}$.
Figure 7.(IV) shows the computation of attention weight $Z$ in column-$i$ chips. Every chip already has the complete duplicate of the column-$i$ $Q$ heads $(2,8,64)$. The key tensor is partitioned horizontally, each chip keeps cached $K$ with shape $(2,64,\ell/4)$ in its local KV-cache ($\ell$ is the current context length). The VEX unit multiplies $Q$ head $(2,8,64)$ with $K$ $(2,64,\ell/4)$ and produces the local attention weight $Z$ $(2,8,\ell/4)$. Because each chip sees only $\ell/4$ tokens, a column-wise all-reduce needed to perform, after which every chip completes the normalisation of its local fragment.
Figure 7.(VI) depicts the output projection and first residual path. After the attention score computation, each of the four chips in a same column now holds the same, flattened attention score for 16 heads. The $W_{o}$ matrix is partitioned row-wise across the column group chips, with each column assigned a weight shape of $(1024,2880)$ and each individual chip’s HN array receiving a $(1024,720)$. Each chip generates a partial-$O$ of shape $(1,720)$. These partial-$O$s are combined via one row-wise all-reduce and one column-wise all-gather to yield the final $X_{o}$ of shape (1, 2880) in all chips.
5.3. Feed-Forward Network with MoE
Figure 7.(VII) shows the experts router stage. After the Group Query Attention computation, all chips in the HNLPU contain the complete $(1,2880)$ $X_{o}$ vector. These values are passed through RMSNorm before entering the routing layer. As $W_{\text{rout}}$ only accounts for only about $0.01\%$ of the total weights, we replicated a copy of the router weights on all 16 chips to avoid inter-chip data exchange. After the computation of $X_{\text{norm}}\cdot W_{\text{rout}}$, each chip obtains complete $X_{\text{rout}}$. Next, top-k and softmax operation are performed. The top-k result is used to generate a masked input tensor, $X$, with a shape of $(128,1,2880)$. In this tensor, the values for the top-k experts are set to those of $X_{\text{norm}}$, while all others are set to zero. Additionally, the top-k results are normalized with a softmax to obtain the expert weights.
Figure 7.(VIII) shows the up- and gate-projection stage. Following the top-k masking, each chip processes its masked $X_{\text{mask}}$ tensor, which consists of 8 vectors, each of shape $(1,2880)$. For 128 $X_{\text{mask}}$ $(1,2880)$ in all chips, only k of which are non-zero. The weight of $W_{\text{up}}$ is sliced into 16 tiles, with the shape of $(8,2880,2880)$ for each chip. Then, the local 8 vectors of $X_{\text{mask}}$ are multiplied with the sliced weight $W_{\text{up}}$ of eight experts. The 16 chips produce a total $X_{\text{up}}$ tensor of shape $(128,1,2880)$. Because each chip stores the complete weight matrices for all experts, this step requires no inter-chip data communication. The gate projection follows the same partitioning: $X_{\text{{mask}}}$ is multiplied by the corresponding $W_{\text{gate}}$ slice, yielding a $X_{G}$, $(128,1,2880)$ in total and $(8,1,2880)$ for each chip. Applying the SwiGLU activation to $X_{G}$ and $X_{\text{{up}}}$, and taking the element-wise product, we get the output $X_{t}$ for the subsequent down-projection.
Figure 7.IX shows the down-projection and second residual path. Still, $X_{t}$ in each chip with the shape of (8, 1, 2880)(8,\,1,\,2880) multiplies its down weight slice $W_{\text{down}}$, (8, 2880, 2880)(8,\,2880,\,2880), to produce a partial $X_{\text{down}}$, (8, 1, 2880)(8,\,1,\,2880). Next, the expert weights, which were obtained from the previous stage (VII), are multiplied with the corresponding expert outputs to get the weighted output for each expert. Subsequently, a all-chip all-reduce operation is performed to sum the partial results. The shape of $X_{\text{down}}$ in all chips is from (128, 1, 2880)(128,\,1,\,2880) to (1, 2880)(1,\,2880). The summed $X_{\text{down}}$ is then added to $X_{o}$ to yield the final layer output Y(1, 2880)Y\,(1,\,2880).
@@ 5.4. Pipeline and Scheduling
HNLPU employs a nested pipelining strategy to maximize throughput, consisting of both inter-layer and intra-layer pipelining. Since all layer weights are hardwired onto the metal layers, the wiring weights for each layer have their own corresponding computing resources. HNs of each layer can operate simultaneously, which facilitates the straightforward formation of a pipeline between the model layers. Within a layer, we partition the computation into a six-stage pipeline, as shown in Figure 8. Consequently, HNLPU can process up to 6 $\times$ layer number of batches simultaneously at peak. For the 36-layers LLM, the maximum batch size can theoretically reach 216.
HNLPU employs the batching strategy similar to the continuous batching (Yu et al., 2022). During the prefill phase, there are no dependencies between the input tokens of a sequence. This independence allows for massively parallel processing, with tokens flowing through the pipeline stage-by-stage. Consequently, HNLPU can process up to 216 tokens concurrently during prefill. Conversely, the decode phase is auto-regressive, meaning the generation of each new token is dependent on the completion of the previous one. However, since different sequences are independent, HNLPU can still process up to 216 sequences simultaneously. In summary, HNLPU supports a maximum batch size of 216 sequences. By leveraging continuous batching, the system dynamically schedules new sequences into the batch as soon as slots are freed by completed ones, thereby ensuring high throughput.

### v2

@@ 3.1. Hardwired-Neuron Architecture
Figure 4. Hardwired-Neuron architecture. ❶ A conventional cell-embedding neuron contains 2,880 4b multipliers (16 shown) followed by an 8b×2,880 adder tree, where 2,880 is the hidden size in gpt-oss 120 B; ❷ With ME, Hardwired-Neurons accept 1b serialized inputs (LSB-first), (1) route the inputs multiplying the same weight value to the same region, (2) perform accumulation (POPCNT) on these input signals, (3) perform actual multiplication with 16 multipliers (4 shown), (4) sum the results with a 4b×16 adder tree. Note how ❷ is significantly smaller in area than ❶ by reducing the number of multipliers and the strength of adders.
Step 2: Bit-serialization
If input signals $x_{1},x_{2},\dots,x_{n}$ are in binary format, they could be serialized from the least-significant bit (LSB) to the most-significant bit (MSB) to further simplify the circuit. As shown in Figure 3 (to the right), the single-clock-cycle accumulation could unfold into a multiple-clock-cycle tree of Carry Save Adders (Hwang, 1979), trading off speed for minimized area.
HN is an accumulate-multiply-accumulate unit adopting all of the above-mentioned techniques as shown in Figure 4. The main result of the sophisticated combination of these techniques is that HN lifted the embedding of weight parameters from silicon devices into metal interconnections, as the name Metal-embedding suggests. Conventional neurons are Cell-Embedding (CE, Figure 4 ❶), i.e., weight parameters are written into the silicon device cells composed of different constant-multipliers; HNs are Metal-Embedding (ME, Figure 4 ❷), i.e., the weight parameters are embedded as metal wires. The silicon device cells in HN can be made parameter-independent.
Figure 5. Step-by-step schematic showing how weights are physically embedded in the 3D metal wire topology. HNs are accumulate-multiply-accumulate arithmetic units where each weight parameter is expressed by the source and destination of a metal wire: ❶ $ax_{1}$ by connecting from $x_{1}$ to the blue region; ❷ $ax_{2}$ by connecting from $x_{2}$ to the blue region; ❸ $cx_{3}$ by connecting from $x_{3}$ to the red region; ❹ $cx_{4}$ by connecting from $x_{4}$ to the red region.
Figure 6. Conceptual schematic of Metal-Embedding. ❶ Conventional design embeds weights in the 2D grid of cells. Metal layers are only for physical design (P&R). ❷ Metal-Embedding dramatically increases the density by leveraging the 3D topology structure of metal wires, emulating the axon-dendrite in brains.
Figure 5 shows the weight embedding process through metal wires step-by-step. For each unique value in FP4, there is an accumulator (POPCNT) in the HN. We denote the accumulators corresponding to each unique weight value with different colors (FP4 has 16 unique values, 4 shown in the schematic: blue for $a$; green for $b$; red for $c$; violet for $d$). For each input $x_{i}$, the weight $w$ is embedded as a metal wire connecting input $x_{i}$ to the accumulator for the value $w$. For example, as the first term is $ax_{1}$, a metal wire is built connecting $x_{1}$ to the blue accumulator (Figure 5 ❶). Note how the silicon devices are made parameter-independent: changing the weight from $a$ to $b$ would only change the wire destination from blue to green, while the whole accumulate-multiply-accumulate arithmetic unit is kept unchanged.
To address the imbalance of weight values, the size of accumulators should be made with sufficient slackness. The accumulators could be implemented as multiple slices and be reconfigurable through metal wires. Unused ports on the accumulators are connected with zero inputs (ground).
@@
The key insight here is that, as the metal wire topology is three-dimensional, they could potentially embed information in a much higher density than silicon devices. Current CE methodologies (Figure 6 ❶) fail to recognize metal layers in their architecture design. Metal layers are only considered in physical design (place-and-route, P&R), thus the topology of metal wires does not express any specific information. We view this as a waste of resources. Since silicon devices are the area-limiting factors in the design, routing signals through complex metal wires is virtually free in both area and energy, compared with logic implemented in standard cells. We refer to the novel architecture as Hardwired-Neurons because of its structural similarities with biological neurons. Biological neurons in the brains have complex topology of axon-dendrite interconnections in the first place, only then come synaptic weights.
When adopted alone, the HN architecture increases the density of hardwired LLM by an order of magnitude (from 200+ chips to 16 chips). But more importantly, HN concentrates all the parameter-dependent structures into metal wires, which is a prerequisite step towards introducing the Sea-of-Neurons architecture.
3.2. Sea-of-Neurons Architecture
Up to this point, there are two common concerns to address: 1) The NRE is still high. Even with significantly reduced area of HN, 16 chips still require 16 full mask sets each valued \$ 30 M, that is \$ 480 M. The total NRE still offset most economic interests. 2) What if the weight parameters change? LLM requires at least annual updates to keep competitive, and there would be unforeseen hotfixes. Do we need another \$ 480 M for a Hardwired LPU re-spin?
The key to these concerns is to share and reuse parameter-independent photomasks. As the silicon devices in HN are parameter-independent, we can prefabricate HN arrays with a shared photomask set, then finalize the metal embedding wires with a few additional parameter-dependent photomask layers. By concentrating metal embedding wires into higher level metal layers, the majority of the photomask cost can be saved. The 16 chips could share the same photomask set for the prefabricated HN array, and the photomask set could be reused for future weight update re-spins.
Figure 7. The cross section of chips (Sell et al., 2022). Geometric patterns in semiconductor chips are defined by photomasks. Silicon devices and lower metal layers have finer geometric features and thus require much more expensive photomasks to define them.
Different layers of photomask in the set are valued differently. Generally speaking, higher levels use cheaper photomasks. As shown in Figure 7, silicon devices and lower metal layers have finer feature dimensions and requires high resolution lithographic patterning which is expensive. For example, metal layers at M10–M11 ($\sim$60 nm half-pitch) require Deep Ultraviolet (DUV) single-exposure patterning (193i SE); at M4–M9 ($\sim$40 nm half-pitch), DUV double patterning is required (typically 193i SADP, with some layers modeled as LELE); at M0–M3 ($\sim$20 nm half-pitch), DUV quadruple patterning (193i SAQP) or Extreme Ultraviolet Lithography (EUV SE) is required. FEOL processes making devices and contacts also require expensive EUV or DUV multiple patterning. Top metal layers including M12+ are typically reserved for power delivery networks, clock trees, and I/O peripherals. Therefore, we select M8-M11 (involving 10 layers of DUV photomasks, valued \$ 2.31 M) as the metal-embedding layers.
Figure 8. Sea-of-Neurons Architecture. To the top: All chips share the same prefabricated HN array and 60 layers of photomask, then the embedding metalization (M8-M11) requires additional 10 layers of photomask per chip. To the bottom: The schematic from cross section with opened HN module box, showing devices and local metal wires (M0-M7) inside the HN array that are made homogeneous and mask-sharing. Note all the critical layers requiring EUV are shared.
The integrated circuit design approaches to save photomask costs by semi-custom metalization over a prefabricated array of cells are known as Structured ASIC (Zahiri, 2003), and have emerged throughout history, including gate arrays in the 1970s, sea-of-gates in the 1990s (de Lima and Kinniment, 1995), Altera HardCopy in the 2000s (Mansur, 2008), and Intel eASIC N5X in 2020 (Intel, 2020). As our approach is prefabricating arrays of neurons instead of gates, we refer to it as the Sea-of-Neurons architecture.
Figure 8 illustrates the Sea-of-Neurons architecture. Sea-of-Neurons is a metal-programmable architecture: Weights are programmed into the architecture with M8-M11 metalization over a prefabricated HN array. As 60 out of 70 mask layers are shared (including all critical layers requiring EUV), the photomask cost is significantly reduced from \$ 480 M to \$ 65 M 222\$ 27.69 M (the prefabricated HN array) + \$ 2.31 M (M8-M11 metalization per-chip) $\times$ 16 (number of chips). When the weight parameters change, a re-spin requires only \$ 37 M 333\$ 2.31 M (M8-M11 metalization per-chip) $\times$ 16 (number of chips) as the prefabricated HN array is ready.
The Sea-of-Neurons architecture is compatible with standard ASIC design flow and EDA tools. First, complete the P&R of the HN array module under standard cell constraints within M0-M7. The layout of HN is copied to fill the major part of die area, equipped with all SoC peripherals, power grid, and clock tree. Next, the layout is exported to custom tools which read weight parameters and generate TCL scripts to instruct the connection of metal embedding wires. The generated script is integrated into the overall layout within the P&R EDA tool. The resulting complete design is then subjected to design rule checking (DRC) and layout-versus-schematic (LVS) verification, with detected rule violations resolved through automated local repair. Finally, parasitic extraction and post-layout simulation is conducted to evaluate functional correctness and timing behavior under realistic physical effects. In our experiments, the layouts successfully completed the sign-off checks showing ample routing density margins in both M0-M7 and M8-M11.
@@ 4. Architecture
4.1. Overview
HNLPU is a complete physical implementation of gpt-oss 120 B and its computational process for inference. HNLPU system directly implements Continuous Batching on hardware to fill its pipeline. The hardware receives token IDs and generates token IDs as outputs, operating without a software stack (OS, runtime, library, compilers, frameworks). This pure hardware implementation offers two benefits: 1) It eliminates the heavy software development and maintenance cost. 2) It eliminates the software turbulence and brings more deterministic and predictable system behavior.
HNLPU distributes the weights across 16 chips interconnected via CXL. Besides the HN array, the chips also implement embedding dictionary lookup, Grouped Query Attention (GQA), Mixture-of-Experts (MoE) routing, Root-Mean-Square Normalization (RMSNorm), Swish-Gated Linear Unit (SwiGLU), and logit sampling. A memory subsystem is implemented for the embedding dictionary and the attention buffer (KV Cache), including SRAM and HBM.
@@ 4.2. System Integration
HNLPU system architecture is built upon a 16-module row-column fully-connected fabric. As conceptually illustrated by the logical topology in Figure 9(a), this fabric establishes direct, point-to-point links from each module to all other modules within its row and, simultaneously, to all other modules within its column. This design creates a router-less, low-latency network for efficient collective communication patterns (e.g., All-Reduce). Each compute module is a self-contained unit, equipped with a dedicated HBM for storing the KV Cache and the embedding tables.
@@
    For attention score, query-heads are all-reduced within the same column groups, while key- and value- heads are reduced to the chip-($\ell\bmod 4$), where $\ell$ denotes the sequence length.
@@ 4.3. Single Chip Architecture
Attention Buffer. The on-chip 320 MB Attention Buffer comprises 20,000 banks, each with a 16 KB capacity. Every bank features a 1W1R (one-write, one-read) port configuration with a 32-bit access width. The Attention Buffer primarily functions as a KV Cache for the chip’s assigned attention groups. It offloads excess KV entries to HBM only when the on-chip capacity is exceeded. This buffer also stores activation vectors for residual connections in the FFN blocks.
@@ 5. Execution Dataflow
In this section, we introduce the multi-chip interconnect dataflow of HNLPU, covering the model-to-chip mapping and the computing process of a Transformer block. In Section 5.1, we provide an overview of HNLPU dataflow. Section 5.2 details our pipelining strategy and batching inference scheduling. Detailed description of the dataflow is presented in Appendix A.
@@ 5.1. Dataflow Overview
Figure 10 illustrates the dataflow of HNLPU. Our design is driven by the primary goals of distributing computational load, KV cache memory access, and thermal loads, while minimizing inter-chip data communication.
@@
The FFN implementation, which uses a Mixture-of-Experts (MoE) architecture, assigns eight experts to each chip, allowing for entirely independent computation with no inter-chip communication. A key exception to our partitioning strategy is the router weight matrix, $W_{\text{rout}}$, which is replicated across all chips. This deliberate design trade-off introduces a negligible area overhead—as the router’s weights constitute only about 0.01% of the total model weights—but eliminates the communication latency. Once all transformer blocks have been processed, the final output vector is passed through the unembedding layer, and a new token index is calculated through a sampling operation. Detailed description of the dataflow is presented in Appendix A.
@@ 5.2. Pipeline and Scheduling
HNLPU employs a nested pipelining strategy to maximize throughput, consisting of both inter-layer and intra-layer pipelining. Since all layer weights are hardwired onto the metal layers, the wiring weights for each layer have their own corresponding computing resources. HNs of each layer can operate simultaneously, which facilitates the straightforward formation of a pipeline between the model layers. Within a layer, we partition the computation into a six-stage pipeline, as shown in Figure 11. Consequently, HNLPU can process up to (6 $\times$ #layer) requests simultaneously at peak. For the 36-layers LLM, the maximum batch size can theoretically reach 216.
HNLPU employs the batching strategy similar to Continuous Batching (Yu et al., 2022). During the prefill phase, there are no dependencies between the input tokens of a sequence. This independence allows for massively parallel processing, with tokens flowing through the pipeline stage-by-stage. Consequently, HNLPU can process up to 216 tokens concurrently during prefill. Conversely, the decode phase is auto-regressive, meaning the generation of each new token is dependent on the completion of the previous one. However, since different sequences are independent, HNLPU can still process up to 216 sequences simultaneously. In summary, HNLPU supports a maximum batch size of 216 sequences. By leveraging Continuous Batching, the system dynamically schedules new sequences into the batch as soon as slots are freed by completed ones, thereby ensuring high throughput.